### v1

# Multiple Choice Learning of Low Rank Adapters for Language Modeling

Abstract
We propose LoRA-MCL, a training scheme that extends next-token prediction in language models with a method designed to decode diverse, plausible sentence continuations at inference time. Traditional language modeling is an intrinsically ill-posed problem: given a context, multiple futures may be equally plausible. Our approach leverages Multiple Choice Learning (MCL) and the Winner-Takes-All (WTA) loss to efficiently handle ambiguity through Low-Rank Adaptation (LoRA). We provide a theoretical interpretation of applying Multiple Choice Learning to Language Modeling, assuming the data is generated from a mixture of distributions. To illustrate the proposed approach, we use data sampled from mixtures of Markov chains. We then demonstrate with extensive experiments on real-world visual and audio captioning tasks that our method achieves high diversity and relevance in generated outputs.
**footnotetext: Equal Contribution.
1 Introduction
Predicting what a person will say next or describing the content of an audio or visual scene with text is difficult, if not impossible, to do with perfect accuracy. When the context is not informative enough, external factors may lead to different scenarios or modes of plausible text continuations yang2018breaking ; dieng2020topic ; mei2022diverse . In such ambiguous tasks, the conditional distribution over the space of output sentences given the input context may be multi-modal due to the underlying inherent uncertainty malinin2020uncertainty .
Initially proposed for text processing, transformer-based language models have quickly become a general framework for modeling streams of tokens, which can also represent, for instance, images and audio signals touvron2023llama ; chu2024qwen2 ; liu2023visual . Such models are trained as next-token predictors and allow, by nature, for addressing this uncertainty, by generating plausible output sentences through a wide body of maximum a posteriori (MAP) or sampling-based decoding approaches welleckdecoding . While sampling allows exploration and diversity, it may lead to unreliable responses and requires truncation to avoid unexpected answers, partly due to the overestimation of low-probability tokens zhang2021trading ; hewitt2022truncation . When seeking reliable and expected answers, maximum a posteriori estimation techniques, like Beam Search beamSearch , look for sentences that maximize the model’s likelihood. However, these alternatives have drawbacks as they may lack diversity, be prone to repetition loops keskar2019ctrl , and may sound unnatural holtzmancurious . Some approaches, e.g., Diverse Beam Search (DBS) vijayakumar2018diverse , were therefore proposed to artificially increase the diversity at inference, e.g., through a diversity penalty parameter $\lambda$, to find a tradeoff between generation quality and sample diversity sucontrastive . In contrast with these methods, our approach aims to predict diverse sentences reflecting the ambiguity of the input context.
Multiple Choice Learning (MCL) guzman2012multiple ; lee2016stochastic has emerged as a paradigm for addressing ambiguous tasks. It generally consists of a network with a shared backbone and multiple output heads. During training, it utilizes the winner-takes-all loss for adaptively updating the head that performs the best for each example. This is a competitive training scheme that specializes each model to subsets of the conditional output distribution rupprecht2017learning ; letzelter2024winner . In this paper, we propose incorporating this idea for language model finetuning, leveraging multiple Low-Rank Adapters hu2022lora instead of multiple heads, which may be impractical due to computation requirements and architectural constraints. Our method natively generates diverse and plausible sequences in a single forward pass, aiming to best approximate the conditional output distribution.
Our contributions are as follows:
We propose a new paradigm that adapts MCL for token sequence modeling, with a method, LoRA-MCL, that is particularly suited for efficient finetuning of language models.
We provide a theoretical analysis of our approach, which applies MCL to language modeling. Assuming the sequences are sampled from a mixture of distributions, we explain in particular why LoRA-MCL should enable capturing the modes of the data distribution. We then consider the case of Markov chains, for which we validate our claims with a well-designed toy example.
We undertake an extensive experimental study to validate our approach, on audio and machine vision captioning tasks, demonstrating its wide applicability in challenging real-world scenarios and showing an excellent diversity–quality trade-off. The code will be made publicly available.
2 Problem setup
Let $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in(\mathbb{R}^{d})^{\tau}$ be a sequence of $\tau$ context vector embeddings of dimension $d$. Language modeling aims at learning the law $p(x\,|\,c)=\prod_{t=1}^{T}p(x_{t}\,|\,x_{<t},c)$ using a model $p_{\theta}$ with parameters $\theta$, by minimizing the following negative log-likelihood loss, which is equivalent to the maximum likelihood estimation (MLE):
$\mathcal{L}(\theta)=-\mathbb{E}_{c,x}[\log p_{\theta}(x\,|\,c)]=\mathbb{E}_{c,x}\Big{[}-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)\Big{]}\;,$ (1)
where $x_{<t}$ denotes the sequence of tokens prior to time $t$.
Optimizing (1) is referred to as teacher-forcing williams1989learning , where $p_{\theta}$ is fed with target (instead of predicted) tokens $x_{<t}$ during training sutskever2014sequence . When using a transformer architecture vaswani2017attention , this is implemented with causal attention modules, which allow for computing the conditional distributions in parallel through all time steps within a single forward pass.
During inference, decoding methods welleckdecoding proceed to generating sequences $\hat{x}$ from the trained model $p_{\theta}$ in an auto-regressive fashion. First, they start with a conditional distribution for the first token from the context: $p_{\theta}(x_{1}\,|\,c)$, which allows selecting $\hat{x}_{1}$. Then, for $t\geq 2$ they predict $p_{\theta}(x_{t}\,|\,\hat{x}_{<t},c)$, and select $\hat{x}_{t}$, until reaching either the sequence length limit or an end-of-sentence (EOS) token. The choice of the decoding method used to generate $K$ candidate sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ depends on the purpose of the task, but the general goal is $(i)$ to get highly likely sentences, i.e., ones that maximize $p_{\theta}(\hat{x})$; and $(ii)$ to get sufficiently diverse sentences, as can be measured by $n$-gram similarity ippolito2019comparison .
Although this is a widely adopted paradigm, we show next how this training and decoding pipeline can be improved: instead of artificially generating diversity at inference time, we aim at learning to predict sequences $\hat{x}^{1},\dots,\hat{x}^{K}$ that cover well the modes of the target distribution $p(x\,|\,c)$.
3 Methodology
3.1 Motivation
In language modeling, topic models nigam2000text ; blei2003latent ; dieng2020topic are data-generating processes in which the ground-truth probability distribution of sequences is modeled as a mixture of latent components or topics. For example, the sequence “I am eating …” may have multiple plausible continuations, but the likelihood of each depends heavily on contextual factors such as the speaker’s location, which influences their culinary habits. Each location (or context) can thus be associated with a distinct word distribution. In topic models, data generation proceeds by first sampling a topic $z\in\mathcal{Z}$ for each sentence (usually referred to as a document in the literature), and then sampling words (or $n$-grams) from the distribution associated with that topic.
With this in mind, MLE in (1) may not be suitable yang2018breaking . While MLE is effective for estimating the overall distribution $p(x)$, it does not capture the individual components when $p(x)$ is expressed as a mixture, i.e., $p(x)=\sum_{k}p(z_{k})p(x\,|\,z_{k})$. In such cases, MLE tends to model the aggregate rather than distinguish topic-specific distributions $p(x\,|\,z_{k})$.
3.2 Applying Multiple Choice Learning to language modeling
Our approach is inspired by the multiple choice learning (MCL) literature guzman2012multiple ; lee2016stochastic . We propose the following training scheme, intending to enable the recovery of the different topics $z_{k}$. Instead of a single model, we consider a set of models $(\theta_{1},\dots,\theta_{K})$. Then the objective (1) is replaced by one consisting of iterating between the following two steps:
  1. 1.
    For each training sample $(c,x)$ in the batch $\mathcal{B}$: Compute $p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$, and choose the best model $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$.
  2. 2.
    Compute the winner-takes-all (WTA) loss as:
    $\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})=-\mathbb{E}_{c,x}\Big{[}\max_{k=1,\dots,K}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (2)
    where $\log p(x\,|\,c;\theta_{k})=\sum_{t=1}^{T}\log p(x_{t}\,|\,x_{<t},c;\theta_{k})$, and perform an optimization step.
In practice when using a finite batch size, (2) is approximated with
$\mathcal{L}^{\mathrm{WTA}}(\theta_{1},\dots,\theta_{K})\simeq-\sum_{k=1}^{K}\sum_{\{(x,c)\in\mathcal{B}\;\,|\,\;k^{\star}(x,c)=k\}}\log p(x\,|\,c;\theta_{k})\;,$ (3)
where each hypothesis $k$ is trained only on the subset of examples in which it is the winner, according to the assignment $k^{\star}(x,c)$.
This training procedure, similar to a hard-EM style optimization min2019discrete ; wen2023an , is a competitive training scheme that encourages the different models to explore different areas of the data distribution. However, it is subject to two main issues: First, using $K$ models instead of a single one drastically increases the training time and memory cost, which may be intractable for large language models. Second, the optimization may be subject to collapse, where the same models are chosen as winners through the iterations, leaving the other models untrained. In the next section, we describe how we solve these issues with our approach LoRA-MCL.
3.3 LoRA-MCL method
Multiple choice learning typically alleviates the high training cost issue of $K$ models by training a single model with several heads lee2016stochastic ; lee2017confident . However, we argue that such an approach is not well-suited for fine-tuning language models. First, heads of most language models are quite large (for example in Qwen2-Audio chu2024qwen2 the lm_head has $d\times|\mathcal{V}|=4096\times 156032\simeq 640\mathrm{M}$ params), and standard MCL would not scale easily with the number of heads. Second, the initialization of the heads poses several challenges. Initializing each head with the parameters of the head of the pretrained model requires special care, as the collapse of the predictions is likely given the very similar hypotheses (same parameters). A complete re-initialization of the heads is detrimental to performance, as numerous training iterations would be necessary to reach the same level of knowledge as in the pretrained head. For these reasons, we consider a Low Rank Adapter (LoRA) approach hu2022lora due to its excellent trade-off between performance and computational requirements, as well as its wide adoption in the context of large model fine-tuning.
Let $\theta$ be the parameters of the pretrained base model. At each layer $\ell$ where LoRA is enabled, we use a family of adapters $(A^{k}_{\ell},B^{k}_{\ell})\in\mathbb{R}^{d\times r}\!\times\!\mathbb{R}^{r\times d}$ for $k\in\{1,\dots,K\}$. Let
$\theta_{k}=\theta\cup\left\{(A_{\ell}^{k},B_{\ell}^{k})\;|\;\ell=1,\dots,L\right\}\;,$ (4)
be the set of parameters that are involved in hypothesis $k$, with $L$ being the total number of layers where LoRA is used.
Training in the WTA fashion involves computing $p(x\,|\,c;\theta_{k})$ for $k=1,\dots,K$. To avoid collapse into a single head, we may use a (relaxed) winner-takes-all loss as a training objective:
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{c,x}\Big{[}\sum_{k=1}^{K}q_{k}\log p(x\,|\,c;\theta_{k})\Big{]}\,,$ (5)
where $\{q_{k}\}$ is a set of coefficients that sum to $1$ and that give more weight to the winner index $q_{k^{\star}}$ than to the others, but still provides gradient weight to non-winner heads in order to avoid collapse. In practice, we use $q_{k^{\star}}=1-\varepsilon$ and $q_{k}=\frac{\varepsilon}{K-1}$ for $k\neq k^{\star}$ as in rupprecht2017learning .
3.4 Accelerating LoRA-MCL training with parallelization over the hypotheses
A naive implementation of MCL for winner selection, as in Section 3.2, may require a loop over the $K$ hypotheses in the batch to determine the winner associated with each index of the batch, which would drastically slow down training.
To alleviate this issue, we propose the following methodology. Let $\mathbf{x}\in\mathbb{R}^{b\times\mathcal{T}\times d}$ denote the input of the transformer architecture, where $b$ is the batch size, $\mathcal{T}\triangleq T+\tau$ is the total sequence length, and $d$ is the number of features. We duplicate $\mathbf{x}$, $K$ times along the batch size dimension to get $\mathbf{x}^{\prime}=(\mathbf{x}_{1},\dots,\mathbf{x}_{K})\in\mathbb{R}^{bK\times\mathcal{T}\times d}$.
LoRA units at layer $\ell$ are then computed as:
$\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]\leftarrow\left[\begin{array}[]{cccc}B^{1}_{\ell}A^{1}_{\ell}&0&0&0\\ 0&B^{2}_{\ell}A^{2}_{\ell}&0&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&0&B^{K}_{\ell}A^{K}_{\ell}\end{array}\right]\left[\begin{array}[]{c}\mathbf{x}_{1}\\ \mathbf{x}_{2}\\ \vdots\\ \mathbf{x}_{K}\end{array}\right]+\left[\begin{array}[]{c}f_{\theta}^{\ell}(\mathbf{x}_{1})\\ f_{\theta}^{\ell}(\mathbf{x}_{2})\\ \vdots\\ f_{\theta}^{\ell}(\mathbf{x}_{K})\end{array}\right]\,,$ (6)
where $f_{\theta}^{\ell}$ is the base model (whose parameters remain frozen during training). In practice, this computation is conducted using a Group Convolution operation.111 This can be done by first reshaping $\mathbf{x}^{\prime}$ to shape $(b\times\mathcal{T}\times Kd)$ and applying a nn.Conv1d($Kd$, $Kd$, kernel_size=1, groups=$K$) following the PyTorch layer implementation, reshaping back $\mathbf{x}^{\prime}$ to shape $(Kb\times\mathcal{T}\times d)$ before adding the base model output, and then repeating at the next LoRA unit. While this duplication virtually multiplies the batch size by $K$, the memory overhead remains manageable, assuming that $r\ll d$. The next section is dedicated to the inner properties of the proposed approach.
4 Theoretical Analysis
We justify the use of MCL in the context of language modeling by assuming that the sequence distribution can be expressed as a mixture. We contrast this formulation with standard maximum likelihood estimation (1). In particular, Section 4.1 establishes a connection between our approach and the Expectation-Maximization (EM) algorithm and derives lower and upper bounds on the optimal achievable test loss. We then apply this analysis to the case of Markov chains in Section 4.2, where the dynamic of the method is simulated in a controlled setting in Section 4.3.
4.1 Training dynamics and optimality conditions
For the next-token prediction loss (1), one can show that: $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$, where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p$ mackay2003information . Following the rationale of Section 3.1, let us now assume the data distribution can be written as a mixture. In the case of the WTA loss under this setup, we have the following proposition.
Proposition 1.
Assume the data-generating process is $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$ (Asm. 2) with the number of modes equal to the number of hypotheses. Also assume perfect model expressiveness (Asm. 1), and large enough batch size to approximate the true risk well (Asm. 3). Then:
  • (i)
    The Winner-Takes-All two-step optimization in LoRA-MCL acts as a conditional form of the hard-EM algorithm.
  • (ii)
    Under Assumption 4, and assuming (with one permutation) that $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, $\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{x,c}\left[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,z_{k})\right]$. In this case, we also have:
    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c,z_{k}\big{)}\right]\,,$ (7)
    where $\mathcal{H}(x\,|\,c,z)$ is the conditional entropy given the random variable $z$.
  • (iii)
    We have the following inequalities:
    $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;,$ (8)
    where $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$.
Proof. See Appendix B.∎
$(i)$ in Proposition 1 describes the relationship between LoRA-MCL and the hard-EM algorithm. $(ii)$ provides an expression for the WTA loss as a conditional entropy, $\mathcal{H}(x\,|\,c,z)$, under the assumption of a perfect matching between the hypotheses and the modes. Proposition 1 also establishes in $(iii)$ both a lower and an upper bound on the optimal achievable loss for LoRA-MCL, given in $(a)$ and $(b)$, respectively. Note that the gap between these bounds is $\mathcal{H}(x\,|\,c,z)-\mathrm{min}_{\theta}\,\mathcal{L}(\theta)+\mathrm{log}\,K=-\mathbb{E}_{c}[\mathcal{I}(x,z\,|\,c)]+\mathrm{log}\,K$ where $\mathcal{I}$ denotes the mutual information. Finally, $(c)$ shows in particular that $\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\leq\mathrm{min}_{\theta}\,\mathcal{L}(\theta)$.
4.2 Case of Markov chains
To make the analysis more concrete, we consider the case where the sequence of tokens is generated from Markov chains. Formally, given a finite state space $\mathcal{V}$, a homogeneous Markov chain is defined by an initial distribution $\pi$ over $\mathcal{V}$ and a transition matrix $P\in[0,1]^{\mathcal{V}\times\mathcal{V}}$. A sequence $(x_{1},\dots,x_{T})$ is sampled from the Markov chain if $x_{1}\sim\pi$ and, for each $t\geq 1$, the conditional distribution of $x_{t+1}$ given $x_{t}$ is $x_{t+1}\,|\,x_{t}\sim P_{x_{t},\cdot}$, where $P_{i,j}=\mathbb{P}(x_{t+1}=j\,|\,x_{t}=i)$. In the following, we ignore the initial warm-up phase and we assume that $\pi$ is the stationary distribution of $P$, i.e., such that $\pi P=\pi$. In this case, we will denote $x\sim\mathrm{MC}(P)$.
While the study of the training dynamics of transformers on Markov Chain data has been investigated in previous works edelman2024evolution ; rajaraman2024analysis ; makkuva2024attention ; zekri2024large , our setup instead considers a mixture of Markov chains gupta2016mixtures ; kausik2023learning . Assuming a uniform mixture, the data generating process is $x\sim\frac{1}{K}\sum_{k=1}^{K}\mathrm{MC}(P_{k})$, which is equivalent to:
$z_{k}\sim\mathcal{U}(1,\dots,K),\quad x\,|\,z_{k}\sim\mathrm{MC}(P_{k})\;.$ (9)
When training a language model on such sequences, we have the following Corollary from Prop. 1, where the context $c$ is ignored for simplicity.
Corollary 1.
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have that:
  • (i)
    Whenever the maximum likelihood estimator trained with next-token-prediction (12) reaches its optimal loss $\mathcal{L}(\theta)$, we have
    $\hat{P}(\theta)_{i,j}=\frac{1}{\sum_{s=1}^{K}(\pi_{s})_{i}}\sum_{k=1}^{K}(\pi_{k})_{i}(P_{k})_{i,j}\;,$ (10)
    where $\pi_{k}\in[0,1]^{\mathcal{V}}$ is the stationary distribution of $P_{k}$.
  • (ii)
    The inequality (8) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by weighting the entropy rate of each Markov Chain:
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$ (11)
Proof. See Appendix C. ∎
The entropy $\mathcal{H}(x)$, which is theoretical optimal loss of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
Note that we considered first-order Markov chains in our analysis. We expect the properties generalize to higher orders ($n>1$) by using transition matrices in the form $P\in[0,1]^{\mathcal{V}^{n+1}}$ where $P_{i_{1},\dots,i_{n+1}}=p(x_{t+1}=i_{n+1}\,|\,x_{t}=i_{n},\dots,x_{t-n+1}=i_{1})$. Please refer to Appendix C for further discussions.
4.3 Illustration with synthetic data
Figure 1: Comparison of LoRA-MCL with vanilla maximum likelihood estimation (MLE). The experimental setup is detailed in Section 4.2. (Left) Validation loss over training steps (averaged across three random seeds) for LoRA-MLE (blue) and LoRA-MCL (red). The theoretical optimal MLE loss, approximated via Monte Carlo sampling, is given by the entropy $\mathcal{H}(x)$. The gray shaded region represents the lower and upper bounds of the theoretical optimal MCL loss, as given by $(a)$ and $(b)$ in (8). (Right) Learned transition matrices (top) are compared with the references (bottom). MLE converges approximately toward the weighted average $\bar{P}$ defined as the right-hand side of (10). In contrast, LoRA-MCL successfully recovers the two underlying modes.
To illustrate our approach, let us evaluate our algorithm on a synthetic dataset, for which results are given in Figure 1. We used $\mathcal{V}=\{1,2\}$, and $(P_{1},P_{2})=(P(p_{1},q_{1}),P(p_{2},q_{2}))$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. We sampled data from a mixture of two Markov chains following (9): we generated a sequence by sampling first $k\sim\mathcal{U}\{1,2\}$ and then $P_{k}$. Once $P_{k}$ was set, we sampled uniformly the initial state, then sampled the Markov chain according to the transition matrix until reaching the maximum sequence length ($T=32$ here).
We considered a GPT-2-like architecture radford2019language ; gpt-neo using local-attention suggested by makkuva2024attention to improve convergence on Markov chain data. We then trained the model with $1$ and $2$ hypotheses, using LoRA adapters as described above, $1$ hypothesis corresponding to vanilla maximum likelihood estimation. Training details are provided in Appendix D. Figures 1 and 4 show both the evolution of the losses along training and the predicted transition matrix from the trained models. We see that the optimum is close to being global in this setup. While LoRA-MCL matches can capture the two modes of the mixture, we see that MLE tends to predict the weighted average of the transition matrices given by (10), which is consistent with Prop. 1.
5 Empirical evaluation
We evaluate LoRA-MCL on realistic datasets and large-scale models for audio and image captioning tasks. Predicting a textual description for images or audio signals is an ill-posed problem: from an input image or audio clip, multiple descriptions may be plausible. We argue this is a particularly relevant real-world case of ill-posed problems, where the conditional output distribution is inherently multi-modal. Our goal is to demonstrate that LoRA-MCL provides a competitive approach for capturing these distribution modes when fine-tuning either audio (Sec. 5.2) or vision-language models (Sec. 5.3). We present hereafter the experimental setup. Please refer to Appendix E.2 for further details.
5.1 Experimental Setup
Datasets. We experimented on both Clotho-V2 font2013freesound ; drossos2020clotho , and AudioCaps gemmeke2017audio ; kim2019audiocaps datasets for the audio captioning task, while we make use of the TextCaps dataset sidorov2020textcaps for a challenging task for vision language models coined as image captioning with reading comprehension. Table 4 of the Appendix describes the datasets sizes.
Architecture and training details with audio data. We used the instructed version of Qwen2-Audio chu2024qwen2 as the base model, which has ${\sim}8.4$ billion parameters and a vocabulary size $|\mathcal{V}|=156{,}032$. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the layers of the transformers. We used a rank $r$ and scaling factor $\alpha$, with $r=\alpha=8$ unless otherwise stated. We trained for 1 epoch on AudioCaps and 10 epochs on Clotho with a validation loss computed at the end of each epoch.
Architecture and training details with visual data. We used LLaVA 1.6 liu2023visual with Vicuna-7B zheng2023judging for the LLM, as the base model which features ${\sim}7.1$ billion parameters and a vocabulary size $|\mathcal{V}|=32{,}000$. We used the official codebase for the implementation. Following standard practices zhou2024empirical , we applied LoRA adapters only for the LLM decoder. The adapters were applied to $Q,K,V$, upside and downside projections as in Qwen2-Audio, and we used $r=\frac{\alpha}{4}=8$ unless otherwise stated. Training was done over 1 epoch (without validation data), and the validation set of TextCaps was used for evaluation.
Metrics. We evaluate the approach using various quality metrics. First, we compute the test-loss (or perplexity), which evaluates the likelihood of the reference captions under the trained model. Second, we use natural language generation quality metrics (BLEU-$n$, ROUGE, METEOR) papineni2002bleu ; lin2004rouge ; banerjee2005meteor based on $n$-gram matching between predicted and target sentences. We also leverage the latter proposed CIDEr, SPICE and SPIDEr vedantam2015cider ; anderson2016spice ; liu2017improved metrics, which show higher correlation with human evaluation for captioning tasks. Finally, we use Sentence-BERT (sBERT) reimers2019sentence ; zhou2022can for embedding comparison. We consider sentence-based oracle evaluation for these metrics (see e.g., (lee2016stochastic, , Sec. 4.3) or labbe2022my ). To evaluate the diversity of the generated captions, we consider: Div-$n$, which compares the number of distinct $n$-grams against the number of $n$-grams for a set of generated captions as in mei2022diverse and Self-BLEU (mBLEU-4), which computes BLEU similarity scores between generated captions zhu2018texygen .
Baselines. To validate our approach, we compare it against a maximum likelihood baseline (LoRA-MLE) trained using (1), under the same conditions as the ones considered for our multi-hypothesis model. Specifically, both models use the same LoRA configuration, the same number of trainable parameters (the LoRA rank for the baseline is $K\times$ larger to this end), and the same number of iterations. At inference time, for each decoding method applied to the baseline that returns $K$ sentences, we decode the same number of candidates with LoRA-MCL. When evaluating Maximum A Posteriori methods –greedy, beam search beamSearch and diverse beam search vijayakumar2018diverse –, we ensure a consistent computational budget by aligning the number of forward passes. For example, if the MLE baseline uses a beam size of $B$, our model uses a beam size of $\frac{B}{K}$ per hypothesis.
5.2 Audio captioning
Figure 2: Quality vs. Diversity on Audio Captioning with $5$ candidates. Quality measured by SPIDEr ()($\uparrow$)( ↑ ) and Diversity by mBLEU-4 ()($\downarrow$)( ↓ ). Marker shape stands for the method, and size is proportional to the number of forwards performed per example at inference time. Note that LoRA-MLE was trained with two rank values, $r=8$ and $r=8K$, for fair comparison in terms of parameter count. Color corresponds to the decoding method (beam search or diverse beam search). Values of $\varepsilon$ and $\lambda$ for LoRA-MCL and diverse beam search, resp., are indicated in the plot, with the color shade proportional to the corresponding parameter value to better distinguish the markers.
Quality vs. diversity trade-off. Quality–diversity performance is shown in Figure 2. While a well-chosen value of $\lambda$ allows Diverse Beam Search applied to LoRA-MLE to be competitive, LoRA-MCL (circles) achieves the best trade-off between quality and diversity, appearing in the top-left corner of the plot. Although increasing the beam size generally improves performance for standard beam search, we observe that increasing the beam size within each group negatively impacts Diverse Beam Search. In theory, setting $\varepsilon=0$ should maximize both quality and diversity. However, using a small positive value of $\varepsilon$ ensures all heads are trained, albeit at the cost of reduced diversity and potentially quality when $\varepsilon$ is too large. Full results are given in Tables 5 and 6 of the Appendix.
Effect of the number of hypotheses. We include in Figures 6 and 5 (given in the appendix) the same plot as in Figure 2, but using 3 and 4 candidates, respectively. These results further reinforce the findings obtained with 5 hypotheses. In Table 5.2, we report the evolution of the test negative log-likelihood (NLL) loss as a function of the number of hypotheses, showing a monotonically decreasing trend. This provides additional evidence for Proposition 1, and shows that LoRA-MCL improves its coverage of the data distribution modes as $K$ increases.
Table 1: Test Loss ()($\downarrow$)( ↓ ) as a function of the number of hypotheses $K$. LoRA-MCL is trained with $\varepsilon=0.05$ and $r=8$.
Training $K$ AudioCaps Clotho
LoRA-MLE ($r=8$) 1 2.152 2.811
LoRA-MLE ($r=40$) 1 2.124 2.902
LoRA-MCL 2 2.071 2.700
LoRA-MCL 3 2.021 2.665
LoRA-MCL 5 1.927 2.629
LoRA-MCL 8 1.869 2.592
Table 2: Quality and Diversity Evaluation on TextCaps with $3$ candidates. Best scores are in bold, second-best are underlined. For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ). LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=4\times r=96$, ensuring the same dynamics and number of parameters across models. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
Training Decoding Beam mBLEU-4 BLEU-4 METEOR sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE Beam Search 3 0.688 0.318 0.315 0.670 1.517 0.244 0.873
LoRA-MLE Beam Search 6 0.786 0.338 0.326 0.671 1.557 0.246 0.895
LoRA-MLE DBS ($\lambda=0.8$) 3 0.437 0.349 0.327 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.416 0.348 0.326 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=1.0$) 3 0.425 0.357 0.327 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.8$) 6 0.671 0.341 0.328 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.666 0.340 0.328 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=1.0$) 6 0.671 0.341 0.328 0.680 1.584 0.251 0.908
LoRA-MCL Beam Search 1 0.520 0.344 0.330 0.690 1.674 0.255 0.955
LoRA-MCL Beam Search 2 0.490 0.360 0.333 0.687 1.627 0.258 0.932
5.3 Image description with reading comprehension
5.3.1 Quality and Diversity Evaluation
Similarly to Figure 2, Table 2 confirms the trends observed in the audio captioning task. At an equal number of forward passes, LoRA-MCL outperforms LoRA-MLE, even using DBS decoding with a $\lambda$ diversity parameter specifically optimized for the task (SPIDEr of $0.955$ against $0.915$). However, we noticed that the DBS decoding of LoRA-MLE tends to generate more diverse outputs than the greedy decoding of LoRA-MCL (mBLEU of $0.416$ against $0.520$). Interestingly, increasing the number of beams in each group decreases diversity and does not improve the performance of LoRA-MLE, while with LoRA-MCL, increasing the beam size (from 1 to 2) not only slightly increases the performance but also improves the diversity of the generations, indicating promising further inference-time scaling. Additional results are provided in Appendix (Table 9).
5.3.2 Observing hypothesis specialization in bilingual image description
To highlight the behavior of LoRA-MCL in a realistic case where one can control the modes of the data-generating process, we simulated an artificial bi-modal distribution of the dataset, similarly to the setup of the toy experiment of Section 4.3. We did so by translating half of the captions of the data from English to French (using T5-small T5 ), while keeping the prompts in English.
We trained and evaluated a 2-hypothesis model in the LoRA-MCL setup, as well as a baseline trained in LoRA-MLE. We observed a clear specialization of each hypothesis towards a given language (i.e., one hypothesis learned French and the other English): at test-time, the winning head is the first one in $\sim 89\%$ of the French captions and the second one in $\sim 97\%$ of the English captions. Table 3 shows performance and diversity scores of both LoRA-MLE and LoRA-MCL on our synthetic test set. LoRA-MCL is decoded greedily while LoRA-MLE uses DBS with the $\lambda$ value maximizing its performance (0.8). Although the overall performances are similar, LoRA-MCL generates significantly more diverse captions (mBLEU-4 of 0.027 and 0.029 against 0.138 and 0.126 on the French and English subparts, respectively), and outperforms LoRA-MLE in the French part of the test set (SPIDEr of 0.464 against 0.411).
These results relate to the synthetic data experiments of Sec. 4.2: in the presence of two modes, LoRA-MLE learns an average of them. Here, we believe this is more shifted toward English as the original LLaVA model performs better in English than in French. In contrast, LoRA-MCL is expected to learn the two modes separately. We believe this is what explains the gain in performance in French, at the cost of a slight reduction in English performance.
Fig. 3 qualitatively illustrates the behavior of the models: LoRA-MLE learns a weighted average of the two modes, shifted towards English, and it is sometimes not able to output French captions within the 2 candidates as in the book example. We found LoRA-MLE to be particularly prone to errors when outputting French sentences: for instance, in the second generation of examples with an image of a bottle, LoRA-MLE enters a repetition loop. On the other hand, LoRA-MCL is less affected by those artifacts on French sentences, and successfully captures the two modes of the distribution, benefiting from the specialization of each hypothesis.
Table 3: SPIDEr ()($\uparrow$)( ↑ ) & mBLEU-4 ()($\downarrow$)( ↓ ) on different parts of synthetic test set.
Test subset Training SPIDEr mBLEU-4
French LoRA-MLE 0.411 0.138
LoRA-MCL 0.464 0.027
English LoRA-MLE 0.756 0.126
LoRA-MCL 0.722 0.029
LoRA-MLE.
{A bottle of Cerveza is on a table.}
{Une bouteille de vin de cidre de cidre de cidre […]}
LoRA-MCL.
{A bottle of beer with a label that says "Sel Maguet"}
{Une bouteille de vin est étiquetée avec le mot « Maguay ».}
LoRA-MLE.
{A book titled Papa Told Me is being held by a person.}
{A book called Papa told me is being held by a person.}
Lora-MCL.
{A book titled Papa Told Me is being held by a person}
{Un livre papier intitulé Papa Told Me.}
Figure 3: Observing specialization in bilingual image description. Quantitative (Left) and Qualitative (Right) analysis for LoRA-MLE and LoRA-MCL in the setup of Section 5.3.2.
6 Related Work
Multiple Choice Learning to predict diverse and plausible outputs. Multiple Choice Learning guzman2012multiple ; lee2016stochastic is a training paradigm that minimizes the oracle or winner-takes-all loss across a set of models, encouraging specialization rupprecht2017learning ; letzelter2024winner . While it has been noted that the collapse issue needs to be addressed with care rupprecht2017learning ; amcl , MCL has demonstrated broad applicability across machine learning tasks lee2017confident ; seo2020trajectory ; garcia2021distillation , typically using a shared backbone and multiple projection heads. However, using multiple heads may be impractical in large language models. Mixture-of-Experts jacobs1991adaptive ; shazeer2017outrageously can be an effective strategy to manage computational costs in this case. However, as primarily thought for model scalability and not diversity, only a small subset of experts are active at each forward pass via routing, making them incompatible with MCL, while still suffering from knowledge redundancy in the experts jiang2024mixtral . To the best of our knowledge, we are the first to adapt MCL to the setup of language modeling with next-token prediction, specifically, using internal models instantiated via multiple LoRA hu2022lora .
Generating Multiple Outputs with Language Models. Language models are commonly trained via next-token prediction, framed as Maximum Likelihood Estimation. This is arguably the most popular method for training large-scale language models shlegerislanguage ; radford2019language ; touvron2023llama , including those that take audio or images as input chu2024qwen2 ; liu2023visual , with much of its success attributed to tokenization makkuva2024attention ; rajaraman2024analysis . Generating diverse and plausible sequences at inference remains challenging: $(i)$ Sampling methods holtzmancurious ; fan2018hierarchical ; meister2023locally may be unreliable depending on the chosen parameters; $(ii)$ Exact MAP decoding is intractable due to the exponential search space eikema2020map ; $(iii)$ Strategies like Beam Search often yield repetitive or overly coherent outputs fan2018hierarchical ; holtzmancurious ; keskar2019ctrl . Diverse Beam Search mei2022diverse or contrastive decoding sucontrastive inject diversity through test-time parameters (e.g., penalty $\lambda$), but in contrast, our method infers diversity from the input’s inherent ambiguity.
Diversity in Audio and Visual Captioning. Audio drossos2017automated ; mei2021audio ; mei2022automated and image captioning aneja2018convolutional ; herdade2019image ; hossain2019comprehensive have traditionally relied on dataset- and task-specific models trained with Maximum Likelihood Estimation (MLE). A key challenge identified in both domains is the lack of diversity in generated captions, prompting the development of other training objectives specifically designed to address this issue mei2022diverse ; xu2022diversity ; xu2024towards ; zhang2024generating ; wang2019describing ; wang2020diversity ; mahajan2020diverse . These methods often require architectural modifications and are typically designed for models trained from scratch. However, recent advances have shifted the field toward general-purpose multimodal large language models chu2024qwen2 ; liu2023visual , which offer broader capabilities but pose new challenges to address the diversity-quality trade-off. In this work, we demonstrate that LoRA-MCL can be effectively applied at the fine-tuning stage of such models.
7 Conclusion
This paper proposes a new paradigm for training language models to generate diverse plausible predictions. This is done with LoRA-MCL, combining the Multiple Choice Learning paradigm and Low-rank Adaptation. We provide a theoretical analysis showing that when the target sequence is written as a mixture, LoRA-MCL is expected to capture the modes of the data distribution. We validate our analysis with experiments on Markov chains, and on the real-world audio and advanced image captioning tasks. Future work includes a careful interpretation of the concepts learned by each model as a function of the ambiguity of the input, e.g., for uncertainty estimation as an alternative to wang2023lora .
Limitations. In LoRA-MCL, setting $\varepsilon$ too high ensures that all hypotheses receive gradients, but it can also overly homogenize the models, thereby reducing diversity. While MCL performs well across a broad range of $\varepsilon$ values, identifying the optimal setting can be challenging. To do so, we plan to incorporate recent annealing MCL techniques, e.g., amcl , which employ a scheduling mechanism for $\varepsilon$ and may help improve the quality of the solution. Note that recent LoRA variants, such as hayou2024lora+ may also be an orthogonal axis of improvement for our method.
Acknowledgments
We thank Mickael Chen for his valuable advice, as well as Gilles Puy for his help in the implementation of the parallelization over the hypotheses. This work was funded by the French Association for Technological Research (ANRT CIFRE contract 2022-1854), and granted access to the HPC resources of IDRIS (allocation 2024-AD011014345) by GENCI.
Appendix A Notations and setup
In the following, $x\triangleq(x_{t})_{t=1}^{T}\in\mathcal{V}^{T}$ be a sequence of $T$ tokens belonging to a finite vocabulary $\mathcal{V}=\{1,\dots,|\mathcal{V}|\}$, and $c\triangleq(c_{t})_{t=1}^{\tau}\in(\mathbb{R}^{d})^{\tau}$ be a sequence of $\tau$ context of embeddings of dimension $d$. In the following $\mathcal{X}=\mathcal{V}^{T}$ and $\mathcal{C}=(\mathbb{R}^{d})^{\tau}$.
Language modeling aims at learning the law $p(x\,|\,c)=\prod_{t=1}^{T}p(x_{t}\,|\,x_{<t},c)$ using a model $p_{\theta}$ with parameters $\theta\in\Theta$, by maximum likelihood estimation, minimizing the negative log-likelihood loss
$\mathcal{L}(\theta)=-\mathbb{E}_{c,x}[\log p_{\theta}(x\,|\,c)]=\mathbb{E}_{c,x}\left[-\sum_{t=1}^{T}\log p_{\theta}(x_{t}\,|\,x_{<t},c)\right]\;,$ (12)
where $x_{<t}$ denotes the sequence of tokens prior to time $t$. In practice, we assume that $p_{\theta}=s_{\eta}\circ f_{\theta}$, where $f_{\theta}(x_{<t},c)\in\mathbb{R}^{\mathcal{V}}$ are the predicted logits and $s_{\eta}:z\in\mathbb{R}^{\mathcal{V}}\mapsto\left(\frac{\mathrm{exp}(z_{j}/\eta)}{\sum_{q=1}^{|\mathcal{V}|}\mathrm{exp}(z_{q}/\eta)}\right)\in[0,1]^{\mathcal{V}}$ is the $\mathrm{softmax}$ operator with temperature $\eta>0$.
In the following, we make the following assumption.
Assumption 1 (Expressiveness).
In the following, we assume that the model $p_{\theta}$ is perfectly expressive. Formally, let $\mathcal{F}_{\theta}\triangleq\{p_{\theta}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$ be the family of conditional distributions realized by the model, where $\mathcal{P}(\mathcal{X})$ is the set of probability distributions on $\mathcal{X}$. We assume the family $\mathcal{F}_{\theta}$ is perfectly expressive, that is $\mathcal{F}_{\theta}=\{p^{\star}:\mathcal{C}\rightarrow\mathcal{P}(\mathcal{X})\}$.
First, note that we have the Proposition 2, a well-known result that is due to the Gibbs inequality (e.g., [42]), for which we provide a proof for completeness.
Proposition 2 ([42]).
Under Assumption 1, for the next-token prediction loss (12), one can show that
$\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)\,,$ (13)
where $\mathcal{H}(x\,|\,c)\triangleq-\mathbb{E}_{c,x}[\mathrm{log}\;p(x\,|\,c)]$ is the entropy of $p(\cdot\,|\,c)$.
Proof.
Let us denote $\mathscr{S}(p)=\{(x,c)\,|\,p(x,c)>0\}$ the support of $p$. We use the convention $p(x,c)\,\mathrm{log}\,p(x,c)=0$ for $(x,c)\in\mathcal{X}\times\mathcal{C}-\mathscr{S}(p)$. Because $\mathrm{log}$ is a concave function, we have under Jensen’s inequality:
$-\int_{\mathscr{S}(p)}\mathrm{log}\left[\frac{p_{\theta}(x\,|\,c)}{p(x,c)}\right]p(x,c)\;\mathrm{d}x\mathrm{d}c\geq-\mathrm{log}\left(\int_{\mathscr{S}(p)}\frac{p_{\theta}(x\,|\,c)}{p(x,c)}p(x,c)\;\mathrm{d}x\mathrm{d}c\right)\geq 0\;.$ (14)
However, because of the convention, the left-hand side of (14) is also equal to the integral over $\mathcal{X}\times\mathcal{C}$. This shows:
$-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p_{\theta}(x\,|\,c)\;\mathrm{d}x\mathrm{d}c\geq-\int_{\mathcal{X}\times\mathcal{C}}p(x,c)\;\mathrm{log}\;p(x\,|\,c)\;\mathrm{d}x\mathrm{d}c=\mathcal{H}(x\,|\,c)\;,$
where the equality is reached for parameter $\theta$ such that $p_{\theta}=p$, whose existence is guaranteed by Assumption 1. ∎
In the following, we denote the Kullback–Leibler divergence between two distributions $\alpha$ and $\beta$ as $\mathrm{KL}(\alpha\,\|\,\beta)\triangleq\int_{\mathscr{S}(\beta)}\alpha(x)\mathrm{log}\frac{\alpha(x)}{\beta(x)}\mathrm{d}x$, where $\mathscr{S}(\beta)\triangleq\{x\in\mathcal{X}\,|\,\beta(x)>0\}$. Note that we have the equality:
$-\mathbb{E}_{\alpha}[\mathrm{log}\,\beta(x)]=\mathrm{KL}(\alpha\,\|\,\beta)+\mathcal{H}(\alpha)\;,$ (15)
where the left-hand side is usually referred as the Cross-Entropy, and $\mathcal{H}(\alpha)\triangleq-\mathbb{E}_{\alpha}[\mathrm{log}\,\alpha(x)]$ is the entropy of $\alpha$. When the context is clear, we will also write the entropy of a distribution $\alpha$ as $\mathcal{H}(x)$ where $x\sim\alpha$.
Appendix B Proof of Proposition 1
Let us now consider the following assumptions.
Assumption 2 (Mixture of latent processes).
The data-generating process writes in form $p(x\,|\,c)=\sum_{k=1}^{K}p(z_{k}\,|\,c)\;p(x\,|\,z_{k},c)$. The Mixture is said to be uniform if $\forall k,p(z_{k}\,|\,c)=\frac{1}{K}$.
Assumption 3 (Minimization of the true risk).
The batch size is large enough so that the minimization of the empirical risk comes down to minimizing the true risk (12).
Remark 1.
Under Assumptions 2 and 3, the optimal reachable loss by maximum likelihood estimation is $\min_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{c}\left[\mathcal{H}(x\,|\,c)\right]$, where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k},c)$.
Assumption 4 (Disjoint components).
This assumption states that $p(x\,|\,c,z_{s})=0$ when $p(x\,|\,c,z_{k})>0$, for $s\neq k$.
Proposition 3.
Under Assumptions 1, 2, and 3, we have that:
  • (i)
    The Winner-Takes-All two-step optimization in LoRA-MCL acts as a conditional form of the hard-EM algorithm.
  • (ii)
    Under Assumption 4, and assuming (with one permutation) that $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for each $k$, $\mathcal{L}^{\mathrm{WTA}}(\theta)=-\mathbb{E}_{x,c}\left[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,z_{k})\right]$. In this case, we also have:
    $\mathcal{L}^{\mathrm{WTA}}(\theta)=\mathcal{H}(x\,|\,c,z)\triangleq\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c,z_{k}\big{)}\right]\,,$ (16)
    where $\mathcal{H}(x\,|\,c,z)$ is the conditional entropy given the random variable $z$.
  • (iii)
    We have the following inequalities:
    $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;,$ (17)
    where $\mathrm{min}_{\theta}\;\mathcal{L}(\theta)=\mathcal{H}(x\,|\,c)$.
Proof of (i) First, let us remind that the hard-EM consists of fitting a distribution $p_{\theta}(x,z)$ to observed data $x\sim p(x)$ where $z$ are (unknown) hidden variables. The fitting starts from randomly initialized parameters $\theta$ and latent variables $z$. It consists of repeating the following operations at each iteration $t$ until convergence:
  1. 1.
    (Expectation) $z_{k}^{\star}=\mathrm{argmax}_{k}\;p(x,z_{k};\theta^{(t)})$
  2. 2.
    (Maximization) $\theta^{(t+1)}=\mathrm{argmax}_{\theta}\;p(x,z_{k}^{\star};\theta^{(t})$
Let us define:
$D(\theta,q)\triangleq\int_{\mathcal{X}}\sum_{k=1}^{K}q(k\,|\,x)\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x),\leavevmode\nobreak\ D(\theta)\triangleq\int_{\mathcal{X}}\max_{k=1,\dots,K}\;\mathrm{log}\,p(x,z_{k};\theta)\;\mathrm{d}p(x)\;,$ (18)
where $q$ is a discrete distribution over $\{1,\dots,K\}$ with exactly one non-zero component that controls the assignment of each $x$ to a fixed $k^{\star}$. Let us define $q(\theta)$ as the discrete distribution defined so that $q(k\,|\,x;\theta)\triangleq\mathbf{1}[k=\mathrm{argmax}_{s}\;p(x,z_{s};\theta)]$. Note that then $D(\theta)=D\left(\theta,q(\theta)\right)$.
For the vanilla (or soft) EM algorithm, the complete data log-likelihood $\int_{\mathcal{X}}\mathrm{log}\,p(x;\theta)p(x)\mathrm{d}x$ is expected to increase at each iteration $t$. Similarly, for the hard-EM, we have that the $D(\theta)$ increases at each iteration.
Indeed, the expectation step comes down to computing $q(\theta^{(t)})$. For for the Maximization step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t)})\right)\geq D\left(\theta^{(t)},q(\theta^{(t)})\right)$, by definition. At the next expectation step, we have: $D\left(\theta^{(t+1)},q(\theta^{(t+1)})\right)\geq D\left(\theta^{(t+1)},q(\theta^{(t)})\right)$, because $q(\theta^{(t+1)})$ computes the best assignment given the parameters $\theta^{(t+1)}$. This shows that $D(\theta^{(t+1})\geq D(\theta^{(t)})$.
First note that the main difference compared to the vanilla form of the (hard) EM algorithm, is that the goal is to fit here a conditional distribution $p(x\,|\,c)$ given pair $(c,x)\sim p(c,x)$. Furthermore, step 2 performs a gradient step (of the neural network weights) instead of a full maximization.
Note that under Assumption 2 the complete data log-likelihood writes as $\mathrm{log}\,p(x\,|\,c;\theta)=\mathrm{log}\left[\;\sum_{k=1}^{K}p(x,z_{k}\,|\,c;\theta)\right]$. However, the variables $z_{k}$ are not known in practice. Lora-MCL works by analogy with the Hard-EM algorithm, which consists, in the Expectation step, of picking for each pair $(x;c)$, $k^{\star}(x,c)=\mathrm{argmax}_{k}\;p(x\,|\,c;\theta_{k})$. Indeed, under Assumption 3, each training step of Lora-MCL writes as the optimization of
$\mathcal{L}^{\mathrm{WTA}}(\theta)=-\int_{\mathcal{X}\times\mathcal{C}}\max_{k=1,\dots,K}\mathrm{log}\;p(x\,|\,c;\theta_{k})\mathrm{d}p(c,x)\;,$ (19)
which is expected to decrease at each training iteration. Note that because the loss is bounded from below (by 00), the sequence of real numbers $\{\mathcal{L}^{\mathrm{WTA}}(\theta^{(t)})\}_{t\geq 0}$ is therefore expected to converge.
To conclude, we can view $(\theta_{1},\dots,\theta_{K})$ as the parameters involved in the estimation of the modes of the conditional distribution with $p(x\,|\,c)=\sum_{k=1}^{K}p(x\,|\,c;\theta_{k})p(\theta_{k}\,|\,c)$. Note that the current form of the algorithm does not estimate the weight of each mode $p(\theta_{k}\,|\,c)$, further work could include incorporating scoring heads to estimate $p(\theta_{k}\,|\,c)$ each $k$ as in [35]. ∎
We then expect to be able to recover the distributions (with one permutation) $\{p(x\,|\,c,z_{k})\}$ from estimated $\{p(x\,|\,c;\theta_{k})\}$, assuming identifiability of the data generating mixture, which we expect to be made easier if the components are enough separated (See e.g., [59, Par. 2.5] or [47, Sec. 2.2]).
Proof of (ii) Let us assume that (with one permutation) $p(x\,|\,z_{k},c)=p(x\,|\,c;\theta_{k})$ for $k\in\{1,\dots,K\}$. This is possible thanks to Assumption 1. Let us show that (16).
Let us define:
$\mathcal{X}_{k}(c,\theta)\triangleq\Big{\{}x\in\mathcal{X}\,|\,\mathrm{log}\;p(x\,|\,c,\theta_{k})\geq\mathrm{log}\;p(x\,|\,c,\theta_{s})\;\forall s\in\{1,\dots,K\}\Big{\}}\;.$ (20)
In this case, the WTA loss (2) writes as
$\mathcal{L}^{\mathrm{WTA}}(\theta)$ $=-\int_{\mathcal{C}}\sum_{k=1}^{K}\int_{\mathcal{X}_{k}(c,\theta)}\mathrm{log}\;p(x\,|\,c;\theta_{k})\;p(x\,|\,c)\;\mathrm{d}x\;p(c)\mathrm{d}c$
$=-\int_{\mathcal{C}}\sum_{k=1}^{K}\sum_{s=1}^{K}\int_{\mathcal{X}_{k}(c,\theta)}\mathrm{log}\;p(x\,|\,c;z_{k})\;p(x\,|\,c;z_{s})\;p(z_{s}\,|\,c)\;\mathrm{d}x\;p(c)\;\mathrm{d}c$
[…]
$=-\int_{\mathcal{C}}\sum_{k=1}^{K}\mathcal{H}(x\,|\,c;z_{k})\;p(z_{k}\,|\,c)\;p(c)\mathrm{d}x\;\mathrm{d}c\text{\quad as $\int_{\mathcal{X}_{k}(c,\theta)}p(x\,|\,c;z_{k})\mathrm{d}x=1$.}$
$=\mathbb{E}_{c}\left[\sum_{k=1}^{K}p(z_{k}\,|\,c)\mathcal{H}\big{(}x\,|\,c;z_{k}\big{)}\right]\;.$
Proof of (iii) Let us show that:
$\mathrm{min}_{\theta}\,\mathcal{L}(\theta)-\mathrm{log}\,K\overset{(a)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}^{\mathrm{WTA}}(\theta)\overset{(b)}{\leq}\mathcal{H}(x\,|\,c,z)\overset{(c)}{\leq}\mathrm{min}_{\theta}\,\mathcal{L}(\theta)\;.$
$(a)$: First, we have $\underset{k=1,\dots,K}{\mathrm{max}}\,p(x\,|\,c,z_{k})\leq\sum_{k=1}^{K}p(x\,|\,c,\theta_{k})$. Therefore,
$\mathcal{L}^{\mathrm{WTA}}(\theta)$ $\geq-\mathbb{E}_{x,c}\left[\mathrm{log}\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,c,\theta_{k})\right]-\mathrm{log}\,K$
[…]
$\geq\mathcal{H}\left(x\,|\,c\right)-\mathrm{log}\,K\,.$
Because $\mathrm{min}_{\theta}\,\mathcal{L}(\theta)=\mathcal{H}\left(x\,|\,c\right)$, we have shown that $(a)$ occurs when the KL term vanishes, which is when $p(x\,|\,c)=\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,c,\theta_{k})$.
$(b)$: We have
$\mathcal{L}^{\mathrm{WTA}}(\theta)$ $=-\mathbb{E}_{x}[\underset{k=1,\dots,K}{\mathrm{max}}\mathrm{log}\,p(x\,|\,c,\theta_{k})]$
$=-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\underset{k=1,\dots,K}{\mathrm{max}}\,p(x\,|\,c,\theta_{k})}{p(x\,|\,c,z)}\right]\underbrace{-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\,p(x\,|\,c,z)\right]}_{\mathcal{H}(x\,|\,c,z)}\,.$
Now let us leverage Assumption 1 to choose $\tilde{\theta}_{k}$ such that $p(x\,|\,c,\tilde{\theta}_{k})=p(x\,|\,c,z_{k})$ for each $k\in\{1,\dots,K\}$. In this case, $\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})\geq p(x\,|\,c,z)$ for each $z\in\{z_{1},\dots,z_{K}\}$, and $-\mathbb{E}_{z}\mathbb{E}_{x\,|\,z}\left[\mathrm{log}\frac{\mathrm{max}_{k}\,p(x\,|\,c,\tilde{\theta}_{k})}{p(x\,|\,c,z)}\right]\leq 0$. Then,
$\mathrm{min}_{\theta}\mathcal{L}^{\mathrm{WTA}}(\theta)\leq\mathcal{L}(\tilde{\theta}_{1},\dots,\tilde{\theta}_{K})\leq\mathcal{H}(x\,|\,c,z)\;,$
which proves $(b)$.
Finally $(c)$ can be directly deduced from the inequality $\mathcal{H}(x\,|\,c,z)\leq\mathcal{H}(x\,|\,c)$.∎
Appendix C Proof of Corollary 1
Let us consider the following assumptions.
Assumption 5 (Markov Chain).
We assume that the data-generating process writes as a uniform mixture of Markov chains of order $n\in\mathbb{N}\setminus\{0\}$, that is, for each $t$ and each $k$, $p(x_{t}\,|\,x_{<t},c,z_{k})=p(x_{t}\,|\,x_{t-1},\dots,x_{t-n},c,z_{k})$.
Corollary 2.
As per Assumption 5, let us assume that the data-generating process writes as a uniform mixture of Markov chains of order $n=1$. We denote as $\hat{P}(\theta)\triangleq(p(x_{t+1}=j\,|\,x_{t}=i))_{i,j}$ the predicted transition matrix when using a language model with parameters $\theta$. Under the same assumptions that in Proposition 1, we have:
  • (i)
    Whenever the maximum likelihood estimator trained with next-token-prediction (12) reaches its optimal loss $\mathcal{L}(\theta)$, we have
    $\hat{P}(\theta)_{i,j}=\sum_{k=1}^{K}p(z=z_{k}\,|\,x_{t}=i)(P_{k})_{i,j}=\frac{1}{\sum_{s=1}^{K}(\pi_{s})_{i}}\sum_{k=1}^{K}(\pi_{k})_{i}(P_{k})_{i,j}\;,$
    where $\pi_{k}\in[0,1]^{\mathcal{V}}$ is the stationary distribution of $P_{k}$.
  • (ii)
    The inequality (8) holds in this context, where the conditional entropy $\mathcal{H}(x\,|\,z)$ can be computed by a weighted sum the entropy rate of each of the $K$ Markov Chains:
    $\mathcal{H}(x\,|\,z)=-{\color[rgb]{0,0,0}T}\sum_{k=1}^{K}\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,.$
    The entropy $\mathcal{H}(x)$, which is the lower bound of the MLE baseline, can be computed either exactly for short sequences, or approximated, e.g., through Monte-Carlo integration.
Proof of (i) The Cross entropy of the maximum-likelihood model is optimal whenever $p_{\theta}=p$.
In this case, for each $i,j\in\{1,\dots,|\mathcal{V}|\}$, we have $p_{\theta}(x_{t+1}=j\,|\,x_{t}=i)=p(x_{t+1}=j\,|\,x_{t}=i)$ and
$p_{\theta}(x_{t+1}=j\,|\,x_{t}=i)=\sum_{k=1}^{K}p(z=z_{k}\,|\,x_{t}=i)\,p(x_{t+1}=j\,|\,x_{t}=i,z_{k})\,.$ (21)
Furthermore, by Bayes’ rule, we have:
$p(z=z_{k}\,|\,x_{t}=i)=\frac{p(x_{t}=i\,|\,z=z_{k})p(z=z_{k})}{\sum_{s=1}^{K}p(x_{t}=i\,|\,z=z_{s})p(z=z_{s})}=\frac{(\pi_{k})_{i}}{\sum_{s=1}^{K}(\pi_{s})_{i}}\;,$
because we have assumed a uniform prior over the mixture components, i.e., $p(z=z_{k})=\frac{1}{K}$, and we assumed stationary regime so that $p(x_{t}=i\,|\,z=z_{k})=(\pi_{k})_{i}$. ∎
Proof of (ii) Because we assumed first-order Markov Chains, we have
$\mathcal{H}(x)=\sum_{t=1}^{T}\mathcal{H}(x_{t}\,|\,x_{<t})\,.$
Because we assumed stationary Markov chains, we have that $\mathcal{H}(x_{t}\,|\,x_{t-1})$ doesn’t depend on $t$ and $\mathcal{H}(x_{t}\,|\,x_{t-1})=\mathcal{H}(x_{2}\,|\,x_{1})=-\sum_{i=1}^{|\mathcal{V}|}(\pi_{k})_{i}\Big{[}\sum_{j=1}^{|\mathcal{V}|}(P_{k})_{i,j}\;\mathrm{log}(P_{k})_{i,j}\Big{]}\,$ (see [7] page 66, Theorem 4.2.4). ∎
Note that we considered first-order Markov chains in our analysis. However we expect the properties to generalize to higher orders ($n>1$) by using transition matrices in the form $P\in[0,1]^{\mathcal{V}^{n+1}}$ where
$P_{i_{1},\dots,i_{n+1}}=p(x_{t+1}=i_{n+1}\,|\,x_{t}=i_{n},\dots,x_{t-n+1}=i_{1})\;.$
Appendix D Experimental details of Section 4.3
Figure 4: Comparison of LoRA-MCL with standard maximum likelihood estimation (MLE). The setup mirrors that of Figure 1, but uses different transition matrices. While the overall behavior remains consistent with Figure 1, we observe a distinct transition in the MLE loss. We interpret this as evidence that MLE increasingly incorporates contextual information as training progresses (see also [44]).
The results in Figure 1 were obtained using the setup described in this section. The same illustration with different transition matrices is shown in Figure 4. Note that in both figures, the loss and theoretical quantities are normalized by $T-1$, since the first token is excluded from the computation.
Dataset. We used a uniform mixture of two first-order homogeneous Markov chains with transition matrices $(P_{1},P_{2})=\big{(}P(p_{1},q_{1}),P(p_{2},q_{2})\big{)}$, with $P(p,q)\triangleq\left[\begin{array}[]{cc}1-p&p\\ q&1-q\end{array}\right]\;$ and $p,q\in[0,1]$. For the experiment in Figure 1, we used $(p_{1},q_{1})=(0.2,0.9)$ and $(p_{2},q_{2})=(0.8,0.25)$. Figure 4 uses $(p_{1},q_{1})=(0.7,0.8)$ and $(p_{2},q_{2})=(0.8,0.25)$. The first state of the sequences sampled according to the stationary distribution of the Markov chain given by $\pi_{k}=\frac{1}{p_{k}+q_{k}}(q_{k},p_{k})$ (see e.g., [44]). The sequences have a fixed length of $T=32$.
Architecture. We considered a GPT-2-like architecture [56] using the GPT-Neo implementation [4] using local-attention suggested by [44] to improve convergence on Markov chain data (with window size of $5$). The model has a hidden size of $64$, $2$ layers of transformer blocks with $2$-heads attention. LoRA adapters for the MLE baseline have rank $r=64$, $\alpha=64$, and dropout disabled. To align the numbers of parameters when $K=2$, we used $r=32$ (and $\alpha=32$) for LoRA-MCL. The models have a total of $65{,}536$ trainable parameters over a total of $230{,}912$. Note that aligning the ranks of LoRA-MCL and LoRA-MCL leads to the same conclusions. Weights of the base model were kept frozen (both for LoRA-MLE and LoRA-MCL) to mimic the dynamics on larger language models.
Training details. We used a cosine scheduler with learning rate of $41e-$ 4, weight decay of $31e-$ 3 with AdamW optimizer, with $(\beta_{1},\beta_{2})=(0.9,0.95)$ as in [44]. We used a batch size of $128$, and trained for $500$ iterations, with validation loss computed every $10$ steps. Only the first $450$ iterations are plotted in the Figures 1 and 4. For the LoRA-MCL runs, we trained with vanilla Winner-Takes-All update. Figure 1 (left) shows the mean and standard deviation of the Winner-Takes-All validation loss across three training seeds, for $K=1$ (LoRA-MLE) and $K=2$ (LoRA-MCL).
Theoretical quantities computations To verify the theoretical results, we computed the following quantities:
  • Theoretical Optimal Loss of the MLE model. It is expressed as $\mathcal{H}(x)$ where $x\sim\frac{1}{K}\sum_{k=1}^{K}p(x\,|\,z_{k})$. It can be approximated by Monte-Carlo sampling with samples $(z_{s},x_{s})\sim p(x,z)$ by first sampling modes $z_{s}\sim\mathcal{U}\{1,\dots,K\}$, then $x_{s}\,|\,z_{s}\sim p(x\,|\,z_{s})$, and by computing
    $-\frac{1}{N}\sum_{s=1}^{N}\sum_{t=2}^{T}\mathrm{log}\,(P_{s})_{x_{s,t},x_{s,t+1}}\,.$ (22)
    Here, $(P_{s})_{x_{s,t},x_{s,t+1}}$ denotes the entry of $P_{s}$ at row $x_{s,t}$ and column $x_{s,t+1}$, and $N$ is the total number of samples. We used $N=50{,}000$ here. Note that the term corresponding to $t=1$ in (22) is discarded, as the first token is typically excluded from the loss computation. For the illustration, (22) was normalized by $T-1$, since the first token is discarded.
  • Theoretical Optimal Loss of the MCL model. It is computed by subtracting $\mathrm{log}K$ to the Theoretical optimal loss (22) of the MLE model. To verify $(iii)$ in Proposition 1, we also computed the mixture of entropy rates given by (11), using a normalization factor of $T-1$.
Appendix E Experimental details on the Audio and Image Captioning tasks
E.1 Setup
Dataset statistics are provided in Table 4.
Audio Datasets. We conducted experiments on two audio captioning datasets: Clotho-V2 [14, 10] and AudioCaps [16, 30]. For both datasets, we used the official training, validation, and test splits. Clotho-V2 provides five reference captions per audio clip across all splits, whereas AudioCaps includes a single caption per clip in the training set and five captions per clip in the validation and test sets. During training on Clotho-V2, which is performed on $10$ epochs, at each epoch, one of the five reference captions is sampled uniformly at random for each audio clip.
Vision datasets. We conducted experiments on the TextCaps dataset. Specifically, we used the official training split for training and the official validation split as our test set. Since each image in the dataset is annotated with five different captions, we duplicated each image five times—associating each duplicate with a distinct caption—to ensure that all reference captions are seen during a single training epoch.
Preprocessing of audio data. Following the implementation of [31], for Clotho and AudioCaps raw audio files are resampled from $44.1\,\mathrm{kHz}$ and $32\,\mathrm{kHz}$ respectively to $16\,\mathrm{kHz}$. They are then cropped to a maximum length of 30 seconds for Clotho and 10 seconds for AudioCaps. The data is then fed to the Qwen-2-Audio pipeline, which includes conversion of the raw waveform into a $128$-channel mel-spectrogram, with a window size of $25$ ms and a hop size of $10$ ms.
Preprocessing of image data. Our image pre-processing pipeline follows the recipe of LLaVA (i.e., resizing to $(336,336)$ and normalization using CLIP mean and standard deviation).
For both modalities, we used the HuggingFace transformers [78] and PEFT [46] Python libraries as part of the implementation. Note that for each of the experiments, we set the repetition penalty [29] to $1.1$ for decoding.
Table 4: Statistics of the audio and image captioning datasets. Num. samples includes {train, validation, test} sets, except for AudioCaps, where we used only {train, validation} sets.
Dataset Num. samples Duration (h) Num. Captions Modality
AudioCaps [30] 48,286 134.1 54 K Audio
Clotho [10] 5,929 37.0 30 K Audio
TextCaps [65] 25,119 N/A 126 K Image
E.2 Audio Captioning Experiments
E.2.1 Experimental setup
Architecture and training. We used the Instructed version of Qwen-2-Audio [6] as the base model, which features $\sim 8.4$ billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$ linear projections of the attention modules, and the upside and downside projections of the feedforward blocks, for all the transformer blocks, which include both the audio encoder and language model decoder. We used a rank $r$, with $r=\alpha=8$ unless otherwise stated, with dropout equal to $0.1$. We trained with a batch size of $2$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with minimum learning rate of $10^{-6}$ and maximum learning rate of $10^{-5}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. The validation loss was computed once every epoch.
E.2.2 Additional results
Figure 5: Quality vs. Diversity on Audio Captioning with $4$ candidates.
Figure 6: Quality vs. Diversity on Audio Captioning with $3$ candidates.
Varying the number $K$ of hypotheses. Figures 5 and 6 present the same scatter plots as Figure 2, but with 4 and 3 candidates respectively. The observed trends remain consistent: LoRA-MCL continues to demonstrate an excellent balance between quality and diversity. Notably, the case with $\varepsilon=0$ performs particularly well, which we attribute to improved stability in vanilla WTA when the number of heads is reduced. Note that the detailed results with $K=5$ associated with Figure 2 are given in Tables 5 and 6 for Clotho and AudioCaps, respectively.
Table 5: Results on Clotho-V2 with 5 candidates and MAP Decoding. ‘BS’ and ‘DBS’ stand for beam search and diverse beam search respectively.
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) BS 5 0.298 0.360 0.827 0.651 0.133 0.229 0.443 0.575 0.621 0.175 0.391
LoRA-MLE ($r=40$) BS 5 0.300 0.365 0.819 0.619 0.105 0.213 0.422 0.559 0.556 0.166 0.354
LoRA-MLE ($r=8$) BS 10 0.310 0.386 0.787 0.650 0.139 0.233 0.450 0.576 0.634 0.181 0.400
LoRA-MLE ($r=40$) BS 10 0.314 0.393 0.780 0.623 0.113 0.219 0.426 0.561 0.564 0.170 0.360
LoRA-MLE ($r=8$) BS 25 0.316 0.402 0.764 0.653 0.143 0.236 0.453 0.580 0.639 0.184 0.404
LoRA-MLE ($r=40$) BS 25 0.324 0.411 0.750 0.629 0.121 0.222 0.433 0.563 0.580 0.171 0.368
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.421 0.560 0.524 0.679 0.145 0.241 0.467 0.597 0.674 0.193 0.421
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.422 0.558 0.524 0.652 0.122 0.224 0.445 0.582 0.601 0.176 0.378
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.451 0.606 0.446 0.683 0.149 0.242 0.470 0.600 0.678 0.194 0.423
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.454 0.607 0.444 0.659 0.125 0.227 0.450 0.585 0.612 0.179 0.384
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.466 0.627 0.413 0.685 0.147 0.243 0.471 0.602 0.683 0.195 0.426
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.470 0.628 0.414 0.660 0.124 0.228 0.451 0.588 0.614 0.180 0.386
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.376 0.475 0.722 0.667 0.139 0.238 0.459 0.593 0.656 0.190 0.413
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.377 0.480 0.716 0.642 0.115 0.223 0.439 0.575 0.589 0.179 0.374
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.386 0.491 0.708 0.668 0.139 0.238 0.459 0.595 0.652 0.191 0.412
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.390 0.495 0.704 0.640 0.115 0.223 0.437 0.577 0.585 0.179 0.371
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.392 0.495 0.704 0.664 0.138 0.240 0.455 0.596 0.655 0.192 0.413
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.395 0.498 0.704 0.641 0.112 0.224 0.437 0.578 0.583 0.182 0.372
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.304 0.369 0.813 0.650 0.134 0.229 0.445 0.574 0.625 0.176 0.393
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.305 0.370 0.812 0.619 0.106 0.214 0.423 0.558 0.555 0.166 0.354
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.302 0.366 0.818 0.649 0.133 0.228 0.443 0.574 0.620 0.176 0.391
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.304 0.369 0.814 0.619 0.106 0.214 0.423 0.558 0.556 0.166 0.354
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.302 0.365 0.819 0.648 0.133 0.229 0.444 0.574 0.620 0.177 0.391
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.305 0.369 0.814 0.619 0.105 0.214 0.423 0.558 0.556 0.167 0.355
LoRA-MCL ($\varepsilon=0.0$) BS 1 0.490 0.633 0.389 0.662 0.127 0.231 0.453 0.597 0.643 0.186 0.401
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.464 0.609 0.433 0.672 0.132 0.233 0.459 0.597 0.649 0.185 0.404
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.471 0.623 0.405 0.677 0.141 0.238 0.463 0.600 0.672 0.191 0.419
LoRA-MCL ($\varepsilon=0.0$) BS 2 0.477 0.622 0.407 0.675 0.147 0.237 0.468 0.598 0.670 0.191 0.417
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.449 0.594 0.453 0.687 0.151 0.243 0.471 0.597 0.679 0.193 0.424
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.448 0.598 0.450 0.691 0.157 0.246 0.478 0.600 0.700 0.197 0.436
LoRA-MCL ($\varepsilon=0.0$) BS 5 0.473 0.618 0.420 0.680 0.153 0.242 0.471 0.599 0.689 0.194 0.429
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.443 0.585 0.474 0.688 0.154 0.246 0.475 0.598 0.684 0.197 0.429
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.445 0.592 0.465 0.691 0.165 0.249 0.482 0.601 0.708 0.199 0.442
Table 6: Results for AudioCaps with 5 candidates and MAP Decoding.
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) BS 5 0.238 0.303 0.881 0.601 0.145 0.304 0.490 0.682 0.582 0.239 0.400
LoRA-MLE ($r=40$) BS 5 0.232 0.298 0.888 0.576 0.134 0.296 0.472 0.679 0.510 0.234 0.362
LoRA-MLE ($r=8$) BS 10 0.252 0.331 0.849 0.606 0.156 0.309 0.497 0.686 0.608 0.242 0.414
LoRA-MLE ($r=40$) BS 10 0.246 0.324 0.861 0.581 0.147 0.300 0.479 0.681 0.520 0.238 0.369
LoRA-MLE ($r=8$) BS 25 0.263 0.352 0.825 0.604 0.163 0.310 0.500 0.687 0.616 0.245 0.418
LoRA-MLE ($r=40$) BS 25 0.258 0.350 0.830 0.583 0.156 0.304 0.483 0.685 0.531 0.240 0.373
LoRA-MLE ($r=8$) DBS ($\lambda=0.2$) 5 0.312 0.454 0.684 0.620 0.166 0.317 0.508 0.697 0.622 0.257 0.421
LoRA-MLE ($r=40$) DBS ($\lambda=0.2$) 5 0.310 0.462 0.676 0.595 0.157 0.314 0.492 0.697 0.534 0.255 0.376
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 5 0.358 0.545 0.531 0.636 0.188 0.327 0.521 0.707 0.664 0.270 0.445
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 5 0.354 0.551 0.528 0.610 0.183 0.324 0.506 0.707 0.577 0.267 0.400
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 5 0.377 0.576 0.483 0.640 0.193 0.330 0.521 0.708 0.672 0.270 0.447
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 5 0.374 0.585 0.473 0.614 0.185 0.324 0.509 0.708 0.588 0.269 0.405
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 5 0.386 0.590 0.460 0.638 0.190 0.329 0.521 0.709 0.668 0.269 0.444
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 5 0.382 0.598 0.452 0.612 0.181 0.323 0.505 0.709 0.581 0.266 0.401
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 10 0.314 0.439 0.777 0.624 0.180 0.318 0.510 0.699 0.636 0.262 0.431
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 10 0.309 0.440 0.780 0.605 0.170 0.315 0.496 0.699 0.558 0.260 0.390
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 10 0.320 0.448 0.770 0.624 0.176 0.318 0.507 0.700 0.626 0.263 0.426
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 10 0.316 0.451 0.771 0.604 0.167 0.313 0.496 0.698 0.556 0.258 0.388
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 10 0.325 0.455 0.764 0.621 0.175 0.317 0.506 0.702 0.632 0.262 0.427
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 10 0.322 0.456 0.769 0.602 0.165 0.313 0.495 0.699 0.553 0.258 0.386
LoRA-MLE ($r=8$) DBS ($\lambda=0.5$) 25 0.247 0.319 0.864 0.604 0.152 0.306 0.493 0.682 0.597 0.243 0.408
LoRA-MLE ($r=40$) DBS ($\lambda=0.5$) 25 0.241 0.315 0.872 0.582 0.141 0.298 0.477 0.683 0.523 0.236 0.370
LoRA-MLE ($r=8$) DBS ($\lambda=0.8$) 25 0.246 0.316 0.867 0.603 0.148 0.305 0.492 0.682 0.592 0.242 0.406
LoRA-MLE ($r=40$) DBS ($\lambda=0.8$) 25 0.240 0.311 0.876 0.581 0.138 0.298 0.475 0.682 0.523 0.235 0.369
LoRA-MLE ($r=8$) DBS ($\lambda=1.0$) 25 0.245 0.314 0.870 0.603 0.150 0.306 0.493 0.682 0.592 0.242 0.406
LoRA-MLE ($r=40$) DBS ($\lambda=1.0$) 25 0.239 0.310 0.878 0.581 0.140 0.298 0.476 0.682 0.525 0.236 0.370
LoRA-MCL ($\varepsilon=0.0$) BS 1 0.496 0.707 0.211 0.649 0.174 0.323 0.518 0.702 0.648 0.255 0.428
LoRA-MCL ($\varepsilon=0.0005$) BS 1 0.450 0.678 0.292 0.654 0.195 0.333 0.530 0.710 0.679 0.272 0.448
LoRA-MCL ($\varepsilon=0.05$) BS 1 0.438 0.663 0.326 0.640 0.198 0.332 0.517 0.710 0.611 0.272 0.414
LoRA-MCL ($\varepsilon=0.0$) BS 2 0.459 0.672 0.233 0.650 0.196 0.330 0.529 0.702 0.679 0.262 0.447
LoRA-MCL ($\varepsilon=0.0005$) BS 2 0.408 0.629 0.348 0.658 0.213 0.336 0.539 0.711 0.716 0.276 0.470
LoRA-MCL ($\varepsilon=0.05$) BS 2 0.397 0.616 0.384 0.632 0.208 0.336 0.524 0.710 0.629 0.274 0.424
LoRA-MCL ($\varepsilon=0.0$) BS 5 0.453 0.657 0.254 0.656 0.211 0.331 0.532 0.702 0.697 0.262 0.456
LoRA-MCL ($\varepsilon=0.0005$) BS 5 0.403 0.615 0.372 0.661 0.229 0.339 0.545 0.712 0.744 0.278 0.484
LoRA-MCL ($\varepsilon=0.05$) BS 5 0.392 0.599 0.419 0.640 0.221 0.337 0.530 0.710 0.660 0.274 0.442
Evaluating with Sampling-based decoding. We also report results using sampling-based decoding methods in Tables 7 and 8 for Clotho and AudioCaps. Specifically, we use Top-$k$ sampling with $k=50$, Top-$p$ sampling with $p=0.95$, and Typical sampling with a threshold of $0.95$. For all sampling methods, we apply a repetition penalty of $1.1$, following [29]. In these experiments, the temperature was set to $\eta=1.0$.
We observe that both LoRA-MLE and LoRA-MCL yield significantly higher diversity compared to MAP decoding, albeit at the cost of reduced output quality. This trade-off is evident when comparing results in Tables 5 and 6. While both methods show an overall improvement in quality on the CIDEr and SPIDEr metrics across datasets, performance on other metrics such as METEOR and SPICE remains relatively comparable. In terms of diversity, there does not appear to be a clear trend indicating which method performs best, as the conclusions vary depending on the dataset. These findings suggest the need for further evaluation with different values of epsilon to better understand the quality-diversity trade-off in sampling-based decoding. Additionally, a more in-depth exploration of how the number of generated hypotheses affects sampling quality—and its implications for test-time inference scaling with LoRA-MCL [85]—is left for future work.
Table 7: Results for Clotho with 5 candidates and Sampling-based decoding.
Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) Top-$k$ sampling 0.635 0.814 0.113 0.585 0.073 0.214 0.400 0.590 0.486 0.176 0.315
LoRA-MLE ($r=8$) Typical $p$ sampling 0.645 0.817 0.108 0.589 0.077 0.215 0.402 0.587 0.492 0.177 0.319
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.646 0.818 0.106 0.592 0.074 0.214 0.401 0.589 0.489 0.178 0.318
LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.642 0.816 0.110 0.600 0.079 0.214 0.406 0.588 0.500 0.175 0.322
LoRA-MCL ($\varepsilon=0.0005$) Typical $p$ sampling 0.648 0.815 0.111 0.595 0.082 0.215 0.409 0.584 0.506 0.175 0.327
LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.649 0.817 0.107 0.595 0.079 0.214 0.408 0.586 0.504 0.171 0.325
Table 8: Results for AudioCaps with 5 candidates and Sampling-based decoding.
Training Decoding Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE ($r=8$) Top-$k$ sampling 0.562 0.810 0.134 0.560 0.115 0.299 0.457 0.699 0.442 0.241 0.318
LoRA-MLE ($r=8$) Nucleus (Top-$p$) sampling 0.576 0.810 0.141 0.560 0.121 0.299 0.460 0.699 0.450 0.241 0.325
LoRA-MLE ($r=8$) Typical sampling 0.567 0.809 0.143 0.550 0.115 0.300 0.452 0.697 0.390 0.241 0.294
LoRA-MCL ($\varepsilon=0.0005$) Top-$k$ sampling 0.583 0.830 0.098 0.582 0.115 0.298 0.470 0.700 0.498 0.239 0.344
LoRA-MCL ($\varepsilon=0.0005$) Nucleus (Top-$p$) sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
LoRA-MCL ($\varepsilon=0.0005$) Typical sampling 0.597 0.831 0.098 0.575 0.124 0.298 0.468 0.697 0.497 0.237 0.344
E.3 Image Captioning Experiments
E.3.1 Experimental setup
Table 9: Quality and Diversity Evaluation on TextCaps with $3$ candidates. For each of the presented metrics, higher is better ()($\uparrow$)( ↑ ) except for mBLEU-4 ()($\downarrow$)( ↓ ). LoRA-MCL is trained with $\varepsilon=0.1$, $r=8$ and $\alpha=32$. LoRA-MLE is trained with $r=24$ and $\alpha=96$. For completeness, we also trained LoRA-MLE with $r=8$ and $\alpha=32$ in the rows marked with .
Training Decoding Beam Div1 Div2 mBLEU-4 BLEU-1 BLEU-4 METEOR ROUGE-L sBERT CIDEr-D SPICE SPIDEr
LoRA-MLE Beam Search 3 0.433 0.509 0.688 0.802 0.318 0.315 0.580 0.670 1.517 0.244 0.873
LoRA-MLE Beam Search 3 0.383 0.421 0.833 0.788 0.315 0.317 0.573 0.670 1.517 0.241 0.874
LoRA-MLE Beam Search 6 0.402 0.457 0.786 0.795 0.338 0.326 0.583 0.671 1.557 0.246 0.895
LoRA-MLE Beam Search 6 0.401 0.456 0.784 0.796 0.339 0.327 0.585 0.672 1.572 0.248 0.903
LoRA-MLE DBS ($\lambda=0.5$) 3 0.489 0.600 0.529 0.818 0.345 0.325 0.596 0.684 1.571 0.253 0.902
LoRA-MLE DBS ($\lambda=0.8$) 3 0.530 0.655 0.437 0.824 0.349 0.327 0.601 0.686 1.590 0.251 0.909
LoRA-MLE DBS ($\lambda=1.0$) 3 0.540 0.669 0.416 0.822 0.348 0.326 0.599 0.685 1.586 0.250 0.906
LoRA-MLE DBS ($\lambda=0.5$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=0.8$) 3 0.485 0.597 0.531 0.821 0.346 0.326 0.596 0.685 1.589 0.255 0.912
LoRA-MLE DBS ($\lambda=1.0$) 3 0.535 0.665 0.425 0.827 0.357 0.327 0.601 0.686 1.601 0.252 0.915
LoRA-MLE DBS ($\lambda=0.5$) 6 0.455 0.532 0.694 0.813 0.344 0.328 0.595 0.681 1.580 0.253 0.908
LoRA-MLE DBS ($\lambda=0.8$) 6 0.470 0.549 0.671 0.812 0.341 0.328 0.593 0.681 1.573 0.251 0.903
LoRA-MLE DBS ($\lambda=1.0$) 6 0.474 0.553 0.666 0.812 0.340 0.328 0.592 0.680 1.577 0.250 0.904
LoRA-MLE DBS ($\lambda=0.5$) 6 0.451 0.528 0.697 0.808 0.343 0.329 0.594 0.681 1.586 0.253 0.911
LoRA-MLE DBS ($\lambda=0.8$) 6 0.464 0.542 0.681 0.810 0.340 0.329 0.593 0.680 1.583 0.253 0.909
LoRA-MLE DBS ($\lambda=1.0$) 6 0.473 0.551 0.671 0.810 0.341 0.328 0.592 0.680 1.584 0.251 0.908
LoRA-MCL Beam Search 1 0.499 0.599 0.520 0.828 0.344 0.330 0.597 0.690 1.674 0.255 0.955
LoRA-MCL Beam Search 2 0.507 0.618 0.490 0.824 0.360 0.333 0.604 0.687 1.627 0.258 0.932
We used LLaVA 1.6 as the base model, which features 7.1 billion parameters. We trained using bfloat16 precision. We used LoRA adapters applied to the $Q,K,V$, up and down linear projections of each block of the language model. We used a rank $r$, with $r=8$ and $\alpha=32$ unless otherwise stated, with dropout equal to $0.1$. We trained each model with a batch size of $8$, with AdamW optimizer [40] (with $\beta_{1}=0.9$, and $\beta_{2}=0.98$), weight decay of $0.05$, using a cosine scheduler with maximum learning rate of $10^{-4}$, with a warmup ratio of $0.1$. Gradient clipping is used with a maximum gradient norm of $1.0$. We used 1 epoch for training, where we duplicated the image as many times as the number of its captions, such that the model sees exactly one time each caption.
E.4 Additional results
Table 9 presents the results of LoRA-MCL and LoRA-MLE. Almost all the metrics show the same trend in which LoRA-MCL outperforms LoRA-MLE with Beam Search (BS) and Diverse Beam Search (DBS) in terms of quality, although DBS produces more varied outputs. Depending on the rank ($r=8$ or $r=24$), we found that setting $\lambda=1.0$ or $\lambda=0.8$, respectively, yields the best quality scores for DBS. Similar to our experiments on Audio Captioning, increasing the rank of LoRA-MLE results in a slight degradation in performance but improves diversity. Additionally, increasing the number of beams in BS with LoRA-MLE results in slightly improved performance but reduced diversity. In contrast, increasing the number of beams in DBS with LoRA-MLE (with fixed $\lambda$) leads to declines in both quality and diversity. Interestingly, with LoRA-MCL, increasing the number of beams enhances both performance and diversity here.
E.4.1 Artificial multilingual dataset creation
To evaluate the behaviour of LoRA-MCL under a multi-modal distribution, we simulated an artificial bi-modal dataset by automatically translating half of the captions from English to French using T5-small [57], while keeping the prompts in English. More specifically, we randomly sampled half of the images and translated their five associated captions. All the training parameters are the same as those in the experiments on the original TextCaps dataset, except the learning rate, which we set at $2\times 10^{-5}$ (as the maximum value in the scheduler) in both the LoRA-MLE and LoRA-MCL.
During evaluation, to assess which head is considered as the winner (for the head specialization analysis), we selected the one that maximizes the SPIDEr score over the references of the given sample.
E.5 Computation details
We run the experiments mostly on H100 NVIDIA GPUs with $80$ GB of RAM. Trainings and inferences were launched on a single GPU for the Audio Captioning experiments, and up to $8$ GPUs for the Image Captioning experiments. The total computing resources used for this project, including failed experiments, amount to approximately $20{,}000$ GPU hours.